OpenStudy (anonymous):

Translate the triangle (-3,-2) (-2,-5) (-8,-9) 6 units right and 4 units up. Name the ordered pairs of the translated triangle's vertices

OpenStudy (anonymous):

x co-ordinate of each vertex=-3+6,-2+6,-8+6=? corresponding y co-odinate of each vertex=-2+4,-5+4,-9+4 respectively write in pairs,i.e.,(x,y)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so is this correct (3,2) (4,-1) (-2,-5) <---- answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

WOOHOO